Current position analysis

The following data were taken from the balance sheet of Nilo Company at the end of two recent fiscal years:

Current Year Previous Year
Current assets:
  Cash $663,600 $528,900
  Marketable securities 874,000 624,100
  Accounts and notes receivable (net) 1,076,000 751,000
  Inventories 1,139,300 880,500
  Prepaid expenses 563,500 547,500
  Total current assets $4,316,400 $3,332,000
Current liabilities:
  Accounts and notes payable (short-term) $957,000 $793,400
  Accrued liabilities 363,000 396,600
  Total current liabilities $1,320,000 $1,190,000

This information has been collected in the Microsoft Excel Online file. Open the spreadsheet, perform the required analysis, and input your answers in the questions below.

Open spreadsheet

  1. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round the ratios to one decimal place and the working capital to the nearest dollar.

    Current Year Previous Year
    1. Working capital $ $
    2. Current ratio
    3. Quick ratio
  2. The liquidity of Nilo has   from the preceding year to the current year. The working capital, current ratio, and quick ratio have all  . Most of these changes are the result of a(n)   in current assets relative to current liabilities.
Answer #1
  1. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round the ratios to one decimal place and the working capital to the nearest dollar.

    Current Year Previous Year
    1. Working capital $4316400-1320000 = 2996400 $3332000-1190000 = 2142000
    2. Current ratio 3.3 2.8
    3. Quick ratio 2613600/1320000 = 2.0 1.6
  2. The liquidity of Nilo has improved from the preceding year to the current year. The working capital, current ratio, and quick ratio have all liquidity ratio . Most of these changes are the result of a(n) increase in current assets relative to current liabilities.